Find f(-2) given that f(x) = 4x^2-2x-25
Answer by ewatrrr(24785)

 
Hi
Find f(-2) given that f(x) = 4x^2-2x-25
substituting -2 for x:
f(-2) = 4(-2)^2 -2(-2) -25 = 16 + 4 - 25 = -5